Consider IPv4, each IP address + a port number gives a unique connection (referred to as Internet socket). If 1/4 of the bits in the port field (16 bits) of the TCP header is fixed for certain reason, how many unique connections can be formed?
Answer #1

Given port field has 16 bits and given 4 bits are fixed so total =16-4=12 bits

Total hosts =2^12 -1

                =4096-1

               =4095 hosts